OpenStudy (anonymous):

Help me please SOMEONE!!! cotx-sin2x=0 How do i find the exact solutions in the interval [0,2pi)

OpenStudy (anonymous):

do you know what sin and cos are at that interval?

OpenStudy (anonymous):

No, just says solve it algebraically in which the solution must be in the given interval

OpenStudy (anonymous):

Ahh! accident

OpenStudy (anonymous):

oh okay

OpenStudy (anonymous):

i didnt mean to delete that.

OpenStudy (anonymous):

sin2x-cotx=0 2sinxcosx-cotx=0 (son2x=2sinxcosx - double angle formula) 2sinxcosx-(cosx/sinx)=0 (cotx=cosx/sinx) sinx(2sinxcosx-(cosx/sinx))=0(sinx) 2sin^2xcosx-cosx=0 (do not cancel out the two cosx's, because you would be losing solutions that way!) cosx(2sin^2x-1)=0 cosx=0 and 2sin^2x-1=0 for cosx=0, you get x=pi/2 and 3pi/2 (in that interval) for 2sin^2x-1=0: 2sin^2x=1 sin^2x=1/2 sinx=plus minus rad (1/2) **rad (1/2) = (rad2)/2 and you know that sin(pi/4)=(rad2)/2 and variations and sin(5pi/4)= - (rad2)/2 and variations so x= pi/4, 3pi/4, 5pi/4, 7pi/4 in the interval

OpenStudy (anonymous):

Replace cot x by (cos x/sin x) and sin 2x by 2sin x*cos x. cos x/sin x - 2sin x*cox x = cos x(1/sin x - 2sin x) = cot x (1 - 2sin^2 x) = = cot x*cos 2x = 0. Solve 2 basic trig equations: a) cot x = 0 --> x = Pi/2, b) cos 2x = 0 --> 2x = Pi/2 and 2x = 3Pi/2 x = Pi/4 x = 3Pi/4